OpenStudy (anonymous):

If given f(x) = –4x – 2 and g(x) = 2x, what is (f · g)(x)? A. (f · g)(x) = –6x – 4 B. (f · g)(x) = –8x2 – 4x C. (f · g)(x) = –6x – 2 D. (f · g)(x) = –16x2 – 4x

OpenStudy (anonymous):

f(x) = –4x – 2 and g(x) = 2x Therefore (f · g)(x) =–4x – 2 -2x= -6x-2 Hence option C is the corect answer. @sirizarry78

OpenStudy (anonymous):

can you help me with a another one please

OpenStudy (anonymous):

@sirizarry78 sure go ahead

OpenStudy (anonymous):

If f(x) = 4x – 7 and g(x) = 2x + 4, evaluate f(x) + g(x) for x = –3. A. –52 B. –21 C. –22 D. –15

OpenStudy (anonymous):

f(x) + g(x) = 4x – 7 +2x + 4 = 6x-3 f(x) + g(x) = 6x-3 substituting x= -3 we find. f(-3) + g(-3) = 6*(-3)-3=-18-3= -21 Hence option B is the correct answer @sirizarry78